OpenStudy (anonymous):

Calculate the volume occupied by 4.25g of CO2 at s.t.p

OpenStudy (anonymous):

it's occupy 2.16 L .

OpenStudy (anonymous):

Moradi thanks for the answer but how did you work it out?

OpenStudy (abhisar):

Any gas at STP occupies 22.4 L. Molar mass of CO2 = 44 g/mol No of moles of CO2 in 4.25 g = 4.25/44=0.096 moles So volume occupied = 22.4*0.096=2.16 L

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

thanks abhisar.

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (anonymous):

I see where I made the mistake. I was multiplying 22.4 by 0.966 instead of 0.096. Got it . Thanks once again.

OpenStudy (abhisar):

Anytime :)

OpenStudy (anonymous):

STEP 1: Find the number of moles of CO2 \[Number of moles of CO _{2} =\frac{ mass }{ molar mass}\] \[= \frac{ 4.25g }{ 44g/mole}\] =0.0966 \[volume of CO _{2} = molar volume x moles = 22.4dm ^{3} \times 0.0966 =2.16dm ^{3}\]

OpenStudy (abhisar):

\(\huge\checkmark\)